For each problem, find the derivative of the function at the given value. Some will require a graphing calculator.


1) y= -(-x+1)^(1/2) at x= -1

2) y= 2sin (2x) at x= - π/2

3) y= x^2 + 4x + 1 at x= -1

4) y= -ln (x) at x = 2




1
Expert's answer
2022-01-30T15:24:25-0500
  1. "y= -(-x+1)^{\frac{1}{2}} \text{ at } x= -1"
"\frac{\mathrm{d}}{\mathrm{d} x}[-\sqrt{1-x}]\\
=-\frac{\mathrm{d}}{\mathrm{d} x}[\sqrt{1-x}]\\
=-\frac{1}{2}(1-x)^{\frac{1}{2}-1} \cdot \frac{\mathrm{d}}{\mathrm{d} x}[1-x]\\
=-\frac{\frac{\mathrm{d}}{\mathrm{d} x}[1]-\frac{\mathrm{d}}{\mathrm{d} x}[x]}{2 \sqrt{1-x}}\\
=-\frac{0-1}{2 \sqrt{1-x}}\\
y'=\frac{1}{2 \sqrt{1-x}}\\[4mm]
\text{ at } x=-1\\
y'=\frac{1}{2 \sqrt{2}}"

2. "y= 2sin (2x) \text{ at } x= - π/2"


"\begin{aligned}
& \frac{\mathrm{d}}{\mathrm{d} x}[2 \sin (2 x)] \\
=& 2 \cdot \frac{\mathrm{d}}{\mathrm{d} x}[\sin (2 x)] \\
=& 2 \cos (2 x) \cdot \frac{\mathrm{d}}{\mathrm{d} x}[2 x] \\
=& 2 \cos (2 x) \cdot 2 \cdot \frac{\mathrm{d}}{\mathrm{d} x}[x] \\
=& 4 \cos (2 x) \cdot 1 \\
y'=& 4 \cos (2 x) 
\end{aligned}
\\[4mm]
\text{ at } x= -\pi/2\\
y' = 4cos(- \pi)\\
y'=4(-1)\\
y'=-4"


3. "y= x^2 + 4x + 1 \text{ at } x= -1"


"\frac{\mathrm{d}}{\mathrm{d} x}\left[x^{2}+4 x+1\right]\\
=\frac{\mathrm{d}}{\mathrm{d} x}\left[x^{2}\right]+4 \cdot \frac{\mathrm{d}}{\mathrm{d} x}[x]+\frac{\mathrm{d}}{\mathrm{d} x}[1]\\
=2 x+4 \cdot 1+0\\
y'=2 x+4 \\[4mm]
\text{ at } x=-1\\
y' = 2(-1)+4\\
y' = 2"

4. "y= -ln (x) \text{ at } x = 2"


"\frac{\mathrm{d}}{\mathrm{d} x}[-\ln (x)]\\
=-\frac{\mathrm{d}}{\mathrm{d} x}[\ln (x)]\\
y'=-\frac{1}{x}\\[4mm]
\text{at } x=2\\
y' = -\frac{1}{2}"